OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). A. y + 3 = -5 (x -2) B. y - 2 = -5 (x - 3) C. y + 3 = -1/5 (x - 2) D. y - 2 = -1/5(x + 3)

OpenStudy (anonymous):

I have no idea how to do this plzzz help

pooja195 (pooja195):

Find the slope first \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

-0.2?

pooja195 (pooja195):

as a fraction -1/5

pooja195 (pooja195):

Once you find the slope plug the info in this formula \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

-3-2= -1/5(2-1) ?

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1) x1=-3 y1=2

pooja195 (pooja195):

leave the x and y alone...

OpenStudy (anonymous):

y-2=-1/5(x-(-3))

pooja195 (pooja195):

and rember that - + - = +

OpenStudy (anonymous):

Would it be D?

pooja195 (pooja195):

D is right :)

pooja195 (pooja195):

great work! :)

OpenStudy (anonymous):

THANKS

pooja195 (pooja195):

;)